# Perturbation theory for gravitational shadows in Kerr-like spacetimes

### v1

@@ I Introduction
Recent observations of black hole shadows by the Event Horizon Telescope collaboration EventHorizonTelescope:2019dse; EventHorizonTelescope:2022wkp; Akiyama:2024zpm have provided significant impetus for the development of novel theoretical approaches to calculating and analyzing gravitational shadows. In these studies, photon spheres and their generalizations play a central role, as they define the boundaries separating particle trajectories that escape to infinity from those captured by the black hole Virbhadra:1999nm; Virbhadra:2002ju; Virbhadra:2008ws; Shoom:2017ril; Gibbons:2016isj; Perlick:2021aok; Cunha:2018acu; Grenzebach:2014fha; Grenzebach:2015oea. These structures are most straightforwardly defined for photons in static spacetimes Claudel:2000yi. In rotating geometries, an analogous role is played by surfaces that host non-planar spherical orbits Teo:2020sey; Kobialko:2020vqf. The concept naturally extends to massive particles and to particles with variable effective mass, such as photons in a dispersive plasma Perlick:2015vta; Perlick:2017fio; Perlick:2023znh; Bezdekova:2022gib; Briozzo:2022mgg; Bogush:2023ojz; Kobialko:2022uzj; Song:2022fdg. As in the vacuum photon case, this framework enables the systematic derivation of analytical expressions for gravitational shadows Kobialko:2023qzo, whose morphology inherently depends on the particle’s energy (or observational frequency). This energy-dependent nature of the shadow allows for the extraction of additional information from observational data Moscibrodzka:2017lcu; Chael:2022meh; Ricarte:2022sxg, thereby establishing the foundation for the shadow spectroscopy Pantig:2025deu.
@@
However, even with an analytical expression available Kobialko:2023qzo, a comprehensive analysis of the gravitational shadow remains a challenging task. This difficulty arises because the shadow boundary is typically parameterized by the radius of the spherical photon orbit, rather than directly by the azimuthal angle on the observer’s celestial sphere. As a result, calculating the standard shadow parameters Johannsen:2013vgc; Cunha:2015yba — such as the horizontal and vertical diameters $D_{X}$ and $D_{Y}$, the center coordinate $X_{C}$, the average radius $\bar{R}$, the deviation from sphericity $\delta C$, and the mean deviation from the Kerr shadow $\delta K$—in a fully analytical manner is non-trivial. In practice, these characteristics usually require numerical computation even when an exact parametric form of the shadow boundary is known. Since performing such numerical calculations over wide ranges of metric parameters can be cumbersome, having reliable analytical approximations is particularly valuable. A series of recent studies, primarily on spherically symmetric spacetimes, has demonstrated the efficacy of a perturbation-theoretic approach Vertogradov:2024dpa; Vertogradov:2024eim; Pantig:2025deu; Pantig:2025eqe; Kobialko:2024zhc. These works have also underscored the necessity for generalizing such methods to stationary, Kerr-like geometries Pantig:2025deu.
@@
We present an analytical approach for calculating key shadow parameters in Benenti-type Papadopoulos:2018nvd; Benenti:1979erw; Demianski:1980mgt stationary metrics using perturbation theory. This method yields key shadow characteristics, such as the diameters and the center position, as simple polynomial expansions accurate to $\sim a^{5}$, where $a$ is the Kerr spin parameter, thereby bypassing the need for repeated numerical integration of cumbersome parametric equations. Benenti metrics represent a broad class of geometries that include many Kerr-like solutions—such as the Kerr-Newman metric—and are characterized by the complete separation of variables in the geodesic equations Papadopoulos:2018nvd; Konoplya:2021slg. Our formalism also incorporates the effects of a plasma medium, a feature of significant relevance for multi-frequency astrophysical observations. The availability of such analytical expressions serves a dual purpose: it provides direct insight into how each additional model parameter influences the shadow’s morphology and significantly simplifies the inverse problem of recovering these parameters from observational data using standard optimization techniques.
@@
The paper is structured as follows. In Section II, we derive a general expression for the gravitational shadow boundary in integrable Benenti-type metrics. Section III presents the derivation of the key shadow parameters using perturbation theory around the Schwarzschild metric. The application of these general formulas to several specific metrics is discussed in Section IV, demonstrating both the high accuracy of the approximation and its utility for metric parameter reconstruction. The main results are summarized in the Conclusion, and an Appendix provides supplementary expressions useful for shadow analysis.
@@ II Setup
This section derives a general expression for the gravitational shadow boundary as perceived by a static asymptotic observer within the framework of Benenti-type metrics Papadopoulos:2018nvd; Papadopoulos:2020kxu; Benenti:1979erw. While expressions of this kind are available in various forms in the literature Grenzebach:2014fha; Grenzebach:2015oea; Konoplya:2021slg; Perlick:2021aok, they provide the necessary foundation for our subsequent perturbation-theoretic analysis. From the outset, we consider the general case of radiation propagating through a dispersive plasma medium Perlick:2015vta; Perlick:2017fio; Perlick:2023znh; Bezdekova:2022gib.
In the geometrical optics approximation, the propagation of radiation in a cold, non-magnetized, and pressureless plasma is governed by the Hamiltonian Perlick:2023znh; Kichenassamy:1985zz
@@
where $p_{\alpha}$ are the photon generalized momenta, $g^{\alpha\beta}$ is the inverse spacetime metric, and $\omega_{p}$ is the plasma frequency. We consider a spacetime metric in the Benenti Papadopoulos:2018nvd; Benenti:1979erw form and assuming a separable plasma distribution Perlick:2017fio; Kobialko:2023qzo. In the coordinates $x^{\alpha}=(r,\theta,t,\phi)$, the metric is given by
@@
where $A$ and $B$ are functions of $r$ and $\theta$, respectively. Indeed, this class of metrics is known to permit complete separation of variables in the Hamilton-Jacobi equations Papadopoulos:2018nvd; Benenti:1979erw.
@@
where $\omega=p_{t}=\partial_{t}S$ and $L=p_{\phi}=\partial_{\phi}S$ correspond to the conserved energy and angular momentum of the photon, $p_{r}=\partial_{r}S$ and $p_{\theta}=\partial_{\theta}S$ are its radial and $\theta$-angular momenta, and $\mathcal{C}$ denotes the Carter constant. We also recall that for an asymptotic observer, $\omega$ represents the observed radiation frequency Perlick:2015vta; Kobialko:2025sls.
It is well known that when a black hole is illuminated by external sources, the boundary of its gravitational shadow is defined by the locus of spherical photon orbits—orbits with a constant radial coordinate $r=\text{const}$ from which light can escape to infinity Perlick:2021aok; Cunha:2018acu; Grenzebach:2014fha; Grenzebach:2015oea; Kobialko:2023qzo. For such spherical orbits, or more generally for photon spheres Virbhadra:1999nm; Claudel:2000yi; Virbhadra:2002ju; Virbhadra:2022ybp, the two conditions $\mathcal{R}=\mathcal{R}^{\prime}=0$ must be satisfied, which are equivalent to
@@
Simultaneously, the condition $\mathcal{J}\geq 0$ defines the latitudinal boundaries of individual photon surfaces that host spherical orbits. The union of all such surfaces constitutes the photon region Grenzebach:2014fha; Grenzebach:2015oea.
@@
where $e^{\alpha}_{t}=v^{\alpha}$, and $e^{\alpha}_{r}$, $e^{\alpha}_{\theta}$, and $e^{\alpha}_{\phi}$ form an orthonormal tetrad Kobialko:2023qzo, $H$ is determined from the condition $p_{t}=-\omega$, where photon 4-momenta reads as $p_{\alpha}=g_{\alpha\beta}\dot{x}^{\beta}$. Here, $\Phi$ and $\Theta$ represent the azimuthal and polar angles on the celestial sphere, respectively, corresponding to the direction from which the radiation arrives. We now introduce the Cartesian coordinates for the stereographic projection of the celestial sphere Kobialko:2023qzo; Bogush:2022hop:
@@
where $a$ Kerr spin parameter arising in the asymptotic limit. It should be noted that this expression, depends on the observer’s frame choice and, for instance, differs from that defined for a Zero Angular Momentum Observer (ZAMO). However, as shown in Kobialko:2022ozq; Kobialko:2023qzo, this difference manifests only as a shift in the shadow’s center, leaving its intrinsic shape unchanged.
@@
Note that this implicit form is particularly convenient for constructing the perturbative expansion. Its internal consistency is verified by substituting the explicit expressions for $R$ and $\rho$ from Refs. Kobialko:2022ozq; Kobialko:2023qzo into Eqs. (12), (13), which yields an identity satisfied for arbitrary metric functions $A$ with correct asymptotic behavior.
@@
The primary characteristics of the shadow are its horizontal and vertical diameters (or sizes) Johannsen:2013vgc; Cunha:2015yba.
@@
The second key characteristic is the position of the shadow’s geometric center. In principle, several methods exist to define this center. One common approach is to compute the centroid (center of mass) of the shadow area via integration Bogush:2022hop.
@@
In this new coordinate system, we can calculate the average shadow radius $\bar{R}_{A,m}$ as
@@
the deviation from perfect sphericity as Johannsen:2013vgc; Cunha:2015yba
@@
For these integrals, all integrand functions must be expressed in terms of the corresponding angular coordinate $\Phi_{A,m}$, or a suitable change of the integration variable must be performed. The set of characteristics $(D_{X},D_{Y},X_{C_{A,m}},\bar{R}_{A,m},\delta C_{A,m},\delta K_{A,m})$ is widely used to provide an invariant description of a gravitational shadow Johannsen:2013vgc; Cunha:2015yba. In the following section, we present fully analytical results for calculating these parameters using perturbation theory.
@@ III Perturbation theory
In this work, we employ the Schwarzschild metric as the unperturbed background and introduce perturbations Vertogradov:2024dpa; Vertogradov:2024eim; Pantig:2025deu; Kobialko:2024zhc corresponding to Kerr rotation, characterized by the spin parameter $a$, and a secondary deformation to the Kerr geometry itself, controlled by a dimensionless parameter $\epsilon$. We assume these perturbations preserve the underlying Benenti-type metric structure Papadopoulos:2018nvd; Benenti:1979erw. The rotation parameter $a$ is treated as the dominant effect relative to the deformation and plasma contributions, adhering to a specific scale hierarchy $\epsilon\lesssim a^{2}/M^{2}$. Consequently, we retain terms up to $\mathcal{O}(a^{5})$ in our expansions, which ensures a cumulative error on the order of $\sim a^{6}$ for the calculated shadow parameters. More precisely, in our perturbative expansion we systematically retain terms
@@
Note that including higher-order terms $\sim a^{6}$ and $\sim\epsilon^{3}$ would significantly complicate the final analytical expressions, whereas truncating the expansion at a lower order would omit information about characteristic asymmetric deformations of the shadow relative to the $Y$-axis. Nevertheless, higher-order terms may become valuable for the numerical problem of reconstructing metric parameters via optimization algorithms Kingma:2014vow.
@@
The next step involves substituting these expansions into Eq. (16) and (17), expanding the resulting expressions as a series up to $\mathcal{O}(a^{5})$, and solving the resulting system of linear equations for the unknown coefficient functions $r^{(i,j)}$ and $R^{(i,j)}$. To manage this algebraically intensive process, we have developed an automated computational algorithm. The resulting general expressions for $R(\Phi)$ are, however, typically cumbersome and not particularly illuminating for direct physical analysis. Therefore, we proceed to calculate directly the key invariant parameters of the shadow Johannsen:2013vgc; Cunha:2015yba: $(D_{X},D_{Y},X_{C_{A,m}},\bar{R}_{A,m},\delta C_{A,m},\delta K_{A,m})$.
@@
$\cos\Phi=\sum^{3}_{i=1}\sum^{8-2i}_{j=1}\Phi_{A}^{(i,j)}\cdot\epsilon^{i-1}\cdot(a/M)^{j-1},$ (50)
Using this relation, we determine the shadow boundary $R_{A}(\Phi_{A})$ in the coordinate system centered on $X_{C_{A}}$. The explicit expression is provided in Appendix, Eq. (87). Integrating $R_{A}(\Phi_{A})$ over a full period then yields the average shadow radius
@@
Thus, all principal characteristics of the gravitational shadow Johannsen:2013vgc; Cunha:2015yba have been derived as relatively simple polynomial expressions. This formulation enables a fully analytical investigation of how various metric parameters influence the shadow’s observable features and help with problem of recovering these parameters from observational data using standard optimization techniques what will be discussed in more detail below.
The perturbative framework may also be worth exploring for an attempt to calculate other key lensing observables, such as image magnifications and distortion parameters Virbhadra:2022iiy; Virbhadra:2024xpk. These quantities are central to Virbhadra’s conjecture on the vanishing sum of image distortions Virbhadra:2022iiy. Testing this in a plasma is notoriously difficult, as it demands extreme ray-tracing precision. Similarly, attempting to apply it to studies of image compactness Virbhadra:2022ybp represents a challenging yet potentially fruitful direction for future work.

### v2

@@ I Introduction
Recent observations of black hole shadows by the Event Horizon Telescope collaboration Akiyama and others (2019, 2022, 2024) have provided significant impetus for the development of novel theoretical approaches to calculating and analyzing gravitational shadows. In these studies, photon spheres and their generalizations play a central role, as they define the boundaries separating particle trajectories that escape to infinity from those captured by the black hole Virbhadra and Ellis (2000, 2002); Virbhadra (2009); Shoom (2017); Gibbons and Warnick (2016); Perlick and Tsupko (2022); Cunha and Herdeiro (2018); Grenzebach et al. (2014, 2015); Stepanian et al. (2021). These structures are most straightforwardly defined for photons in static spacetimes Claudel et al. (2001). In rotating geometries, an analogous role is played by surfaces that host non-planar spherical orbits Teo (2021); Kobialko and Gal’tsov (2020). The concept naturally extends to massive particles and to particles with variable effective mass, such as photons in a dispersive plasma Perlick et al. (2015); Perlick and Tsupko (2017, 2024); Bezdekova et al. (2022); Briozzo et al. (2023); Bogush et al. (2023); Kobialko et al. (2022a); Song and Zhang (2023); Bogush et al. (2024). As in the vacuum photon case, this framework enables the systematic derivation of analytical expressions for gravitational shadows Kobialko et al. (2024), whose morphology inherently depends on the particle’s energy (or observational frequency). This energy-dependent nature of the shadow allows for the extraction of additional information from observational data Moscibrodzka and Gammie (2018); Chael et al. (2023); Ricarte et al. (2023), thereby establishing the foundation for the shadow spectroscopy Pantig and Övgün (2025).
@@
However, even with an analytical expression available Kobialko et al. (2024), a comprehensive analysis of the gravitational shadow remains a challenging task. This difficulty arises because the shadow boundary is typically parameterized by the radius of the spherical photon orbit, rather than directly by the angle on the observer’s celestial sphere. As a result, calculating the standard shadow parameters Johannsen (2013); Cunha et al. (2015) — such as the horizontal and vertical diameters $D_{X}$ and $D_{Y}$, the center coordinate $X_{C}$, the average radius $\bar{R}$, the deviation from sphericity $\delta C$, and the mean deviation from the Kerr shadow $\delta K$—in a fully analytical manner is non-trivial. In practice, these characteristics usually require numerical computation even when an exact parametric form of the shadow boundary is known. Since performing such numerical calculations over wide ranges of metric parameters can be cumbersome, having reliable analytical approximations is particularly valuable. A series of recent studies, primarily on spherically symmetric spacetimes, has demonstrated the efficacy of a perturbation-theoretic approach Vertogradov and Övgün (2024b, a); Pantig and Övgün (2025); Pantig (2025); Kobialko and Gal’tsov (2025); Mishra et al. (2019). These works have also underscored the necessity for generalizing such methods to stationary, Kerr-like geometries Pantig and Övgün (2025).
@@
We present an analytical approach for calculating key shadow parameters in Benenti-Francavilla-type Papadopoulos and Kokkotas (2018); Benenti and Francaviglia (1979); Demianski and Francaviglia (1980) stationary metrics using perturbation theory. This method yields key shadow characteristics, such as the diameters and the center position, as simple polynomial expansions accurate to $\sim a^{5}$, where $a$ is the Kerr spin parameter, thereby bypassing the need for repeated numerical integration of cumbersome parametric equations. Benenti-Francavilla metrics represent a broad class of geometries that include many Kerr-like solutions—such as Kerr-off shell metric Ben Achour et al. (2025) — and are characterized by the complete separation of variables in the geodesic equations Papadopoulos and Kokkotas (2018); Konoplya and Zhidenko (2021). Our formalism also incorporates the effects of a plasma medium, a feature of significant relevance for multi-frequency astrophysical observations. The availability of such analytical expressions serves a dual purpose: it provides direct insight into how each additional model parameter influences the shadow’s morphology and significantly simplifies the inverse problem of recovering these parameters from observational data using standard optimization techniques.
@@
The paper is structured as follows. In Section II, we derive a general expression for the gravitational shadow boundary in integrable Benenti-Francavilla-type metrics. Section III presents the derivation of the key shadow parameters using perturbation theory around the Schwarzschild metric. The application of these general formulas to several specific metrics is discussed in Section IV, demonstrating both the high accuracy of the approximation and its utility for metric parameter reconstruction. The main results are summarized in the Conclusion, and an Appendix provides supplementary expressions useful for shadow analysis.
@@ II Setup
This section derives a general expression for the gravitational shadow boundary as perceived by a static asymptotic observer within the framework of Benenti-Francavilla-type metrics Papadopoulos and Kokkotas (2018, 2021); Benenti and Francaviglia (1979). While expressions of this kind are available in various forms in the literature Grenzebach et al. (2014, 2015); Konoplya and Zhidenko (2021); Perlick and Tsupko (2022); Banerjee et al. (2020); Vincent et al. (2021); Tsukamoto et al. (2014); Tsukamoto (2018), they provide the necessary foundation for our subsequent perturbation-theoretic analysis. From the outset, we consider the general case of radiation propagating through a dispersive plasma medium Perlick et al. (2015); Perlick and Tsupko (2017, 2024); Bezdekova et al. (2022).
In the geometrical optics approximation, the propagation of radiation in a cold, non-magnetized, and pressureless plasma is governed by the Hamiltonian Perlick and Tsupko (2024); Kichenassamy and Krikorian (1985)
@@
where $p_{\alpha}$ are the photon generalized momenta, $g^{\alpha\beta}$ is the inverse spacetime metric, and $\omega_{p}$ is the plasma frequency. We consider a spacetime metric in the Benenti-Francavilla Papadopoulos and Kokkotas (2018); Benenti and Francaviglia (1979) form and assuming a separable plasma distribution Perlick and Tsupko (2017); Kobialko et al. (2024). In the coordinates $x^{\alpha}=(r,\theta,t,\phi)$, the metric and the plasma frequency distribution are given by
@@
where $A$ and $B$ are functions of $r$ and $\theta$, respectively. This class of metrics permits complete separation of variables in the Hamilton-Jacobi equations Papadopoulos and Kokkotas (2018); Benenti and Francaviglia (1979) and is algebraically general Gal’tsov and Kulitskii (2024). It therefore provides a broader parametrization of Kerr-like spacetimes than, for instance, the Kerr-off shell geometry Ben Achour et al. (2025).
@@
where $\omega=p_{t}=\partial_{t}S$ and $L=p_{\phi}=\partial_{\phi}S$ correspond to the conserved energy and angular momentum of the photon, $p_{r}=\partial_{r}S$ and $p_{\theta}=\partial_{\theta}S$ are its radial and $\theta$-angular momenta, and $\mathcal{C}$ denotes the Carter constant. We also recall that for an asymptotic observer, $\omega$ represents the observed radiation frequency Perlick et al. (2015); Kobialko et al. (2025).
It is well known that when a black hole is illuminated by external sources, the boundary of its gravitational shadow is defined by the locus of spherical photon orbits—orbits with a constant radial coordinate $r=\text{const}$ from which light can escape to infinity Perlick and Tsupko (2022); Cunha and Herdeiro (2018); Grenzebach et al. (2014, 2015); Kobialko et al. (2024). For such spherical orbits, or more generally for photon spheres Virbhadra and Ellis (2000); Claudel et al. (2001); Virbhadra and Ellis (2002); Virbhadra (2024a), the two conditions $\mathcal{R}=\mathcal{R}^{\prime}=0$ must be satisfied, which are equivalent to
@@
Simultaneously, the condition $\mathcal{J}\geq 0$ defines the latitudinal boundaries of individual photon surfaces that host spherical orbits. The union of all such surfaces constitutes the photon region Grenzebach et al. (2014, 2015).
@@
where $e^{\alpha}_{t}=v^{\alpha}$, and $e^{\alpha}_{r}$, $e^{\alpha}_{\theta}$, and $e^{\alpha}_{\phi}$ form an orthonormal tetrad Kobialko et al. (2024), $H$ is determined from the condition $p_{t}=-\omega$, where photon 4-momenta reads as $p_{\alpha}=g_{\alpha\beta}\dot{x}^{\beta}$. Here, $\Phi$ and $\Theta$ represent the azimuthal and polar angles on the celestial sphere, respectively, corresponding to the direction from which the radiation arrives. We now introduce the Cartesian coordinates for the stereographic projection of the celestial sphere Kobialko et al. (2024); Bogush et al. (2022):
@@
where $a$ Kerr spin parameter arising in the asymptotic limit. It should be noted that this expression, depends on the observer’s frame choice and, for instance, differs from that defined for a Zero Angular Momentum Observer (ZAMO). However, as shown in Kobialko et al. (2022b, 2024), this difference manifests only as a shift in the chosen location of the shadow’s center, leaving its intrinsic shape invariant.
@@
Note that this implicit form is particularly convenient for constructing the perturbative expansion. Its internal consistency is verified by substituting the explicit expressions for $R$ and $\rho$ from Refs. Kobialko et al. (2022b, 2024) into Eqs. (12), (13), which yields an identity satisfied for arbitrary metric functions $A$ with correct asymptotic behavior.
@@
The primary characteristics of the shadow are its horizontal and vertical diameters (or sizes) Johannsen (2013); Cunha et al. (2015).
@@
The second key characteristic is the position of the shadow’s geometric center. In principle, several methods exist to define this center. One common approach is to compute the centroid (center of mass) of the shadow area via integration Banerjee et al. (2020); Bogush et al. (2022).
@@
In this new coordinate system, we can calculate the average shadow radius $\bar{R}_{A,m}$ as Johannsen (2013); Cunha et al. (2015)
@@
Note that the literature also employs an alternative, area-based definition for the average shadow radius Banerjee et al. (2020); Bambi et al. (2019)
$\bar{R}=\sqrt{\frac{1}{\pi}\int dA}=\sqrt{\frac{1}{2\pi}\int R^{2}\cdot d\Phi},$ (26)
which offers the practical advantage of being calculable directly without shifting to a coordinate system centered on the shadow.
The deviation from perfect sphericity reads as Johannsen (2013); Cunha et al. (2015); Banerjee et al. (2020); Bambi et al. (2019)
@@
For these integrals, all integrand functions must be expressed in terms of the corresponding angular coordinate $\Phi_{A,m}$, or a suitable change of the integration variable must be performed. The set of characteristics $(D_{X},D_{Y},X_{C_{A,m}},\bar{R}_{A,m},\delta C_{A,m},\delta K_{A,m})$ is widely used to provide an invariant description of a gravitational shadow Johannsen (2013); Cunha et al. (2015). In the following section, we present fully analytical results for calculating these parameters using perturbation theory.
@@ III Perturbation theory
In this work, we employ the Schwarzschild metric as the unperturbed background and introduce perturbations Vertogradov and Övgün (2024b, a); Pantig and Övgün (2025); Kobialko and Gal’tsov (2025) corresponding to Kerr rotation, characterized by the spin parameter $a$, and a secondary deformation to the Kerr geometry itself, controlled by a dimensionless parameter $\epsilon$. We assume these perturbations preserve the underlying Benenti-Francavilla-type metric structure Papadopoulos and Kokkotas (2018); Benenti and Francaviglia (1979). The rotation parameter $a$ is treated as the dominant effect relative to the deformation and plasma contributions, adhering to a specific scale hierarchy $\epsilon\lesssim a^{2}/M^{2}$. Consequently, we retain terms up to $\mathcal{O}(a^{5})$ in our expansions, which ensures a cumulative error on the order of $\sim a^{6}$ for the calculated shadow parameters. More precisely, in our perturbative expansion we systematically retain terms
@@
Note that including higher-order terms $\sim a^{6}$ and $\sim\epsilon^{3}$ would significantly complicate the final analytical expressions, whereas truncating the expansion at a lower order would omit information about characteristic asymmetric deformations of the shadow relative to the $Y$-axis. Nevertheless, higher-order terms may become valuable for the numerical problem of reconstructing metric parameters via optimization algorithms Kingma and Ba (2014).
@@
The next step involves substituting these expansions into Eq. (16) and (17), expanding the resulting expressions as a series up to $\mathcal{O}(a^{5})$, and solving the resulting system of linear equations for the unknown coefficient functions $r^{(i,j)}$ and $R^{(i,j)}$. To manage this algebraically intensive process, we have developed an automated computational algorithm. The resulting general expressions for $R(\Phi)$ are, however, typically cumbersome and not particularly illuminating for direct physical analysis. Therefore, we proceed to calculate directly the key invariant parameters of the shadow Johannsen (2013); Cunha et al. (2015): $(D_{X},D_{Y},X_{C_{A,m}},\bar{R}_{A,m},\delta C_{A,m},\delta K_{A,m})$.
@@
$\cos\Phi=\sum^{3}_{i=1}\sum^{8-2i}_{j=1}\Phi_{A}^{(i,j)}\left(\cos\Phi_{A}\right)\cdot\epsilon^{i-1}\cdot(a/M)^{j-1},$ (51)
This expansion is substituted into Eq. (50) to determine all coefficients. Using this result, we obtain the shadow boundary $R_{A}(\Phi_{A})$ in the coordinate system centered on $X_{C_{A}}$. The explicit expression is provided in Appendix, Eq. (90). Integrating $R_{A}(\Phi_{A})$ over a full period then yields the average shadow radius
@@
For the area-based definition (26), we find only a slight adjustment independent of the Kerr metric deformation
$\bar{R}-\bar{R}_{A}=\frac{a^{4}\sin^{4}\bar{\theta}}{576\sqrt{3}M^{3}}+\mathcal{O}(a^{6}).$ (52)
As anticipated, this value is slightly overestimated, owing to the greater weight of shadow boundary regions located farther from the center.
@@
Note that to the order considered, the use of the average radius $\bar{R}$ instead of $\bar{R}_{A}$ leads to the same result. At the same time, using a different center $X_{C_{m}}$ leads to an expected increase in this deviation.
$\delta C_{m}-\delta C_{A}=\frac{\sqrt{2}a^{4}\sin^{4}\bar{\theta}}{2187M^{4}}+\mathcal{O}(a^{6}).$ (54)
@@
Thus, all principal characteristics of the gravitational shadow Johannsen (2013); Cunha et al. (2015) have been derived as relatively simple polynomial expressions. This formulation enables a fully analytical investigation of how various metric parameters influence the shadow’s observable features and help with problem of recovering these parameters from observational data using standard optimization techniques what will be discussed in more detail below.
The perturbative framework may also be worth exploring for an attempt to calculate other key lensing observables, such as image magnifications and distortion parameters Virbhadra (2022, 2024b); Virbhadra and Keeton (2008). These quantities are central to Virbhadra’s conjecture on the vanishing sum of image distortions Virbhadra (2022). Testing this in a plasma is notoriously difficult, as it demands extreme ray-tracing precision. Similarly, attempting to apply it to studies of image compactness Virbhadra (2024a) represents a challenging yet potentially fruitful direction for future work.